OpenStudy (kittykitty04):

Analysis of a 45.0g sample of water revealed a silver concentration of 47.1ppm. Assuming the solution has a density of 1.00g/mL. Calculate the molar concentration of the silver in the solution. (I am so lost with this question. Does it make sense to anyone?) help! :(

OpenStudy (aaronq):

convert 47.1 ppm to grams then to moles... 1g of water = 1 mL then Molarity=moles/L of solution

OpenStudy (kittykitty04):

okay... how do i go from ppm to g? :S

OpenStudy (aaronq):

ppm=(g of silver/g of water) x 1,000,000

OpenStudy (aaronq):

oh wait sort i meant *g of solution at the bottom

OpenStudy (kittykitty04):

okay, so you're saying ppm = (g of water/ g of solution)?

OpenStudy (aaronq):

times a million, yes

OpenStudy (kittykitty04):

sorry, this question confuses me because it says "silver concentration of 47.1 ppm" in the beginning of the question and then it says "calculate the concentration of the lead in the solution" at the end. it totally messes with me! ah! :| So it says "assuming the solution has a density of 1.00g/mL" should I do ( 45.0g water/ 1.00 g/ml) = 45 mL? i'm so lost.

OpenStudy (kittykitty04):

x a million. lol

OpenStudy (kittykitty04):

wait. ugh, this is so frustrating. all that i just said that made no sense.

OpenStudy (aaronq):

haha oh i didn't even notice that.. it's probably a typo..

OpenStudy (aaronq):

47.1= (g of silver/45 mL )x1,000,000

OpenStudy (aaronq):

45 mL can from mass of water and density 1 g/ml= 45 g/V V= 45 mL

OpenStudy (aaronq):

45 came*

OpenStudy (kittykitty04):

alright. i think i got it, thanks aaron!

OpenStudy (aaronq):

(Y) good stuff !